A work is done by 66 workers not all of them have the same capacity to work. Every day exactly 2 workers, do the work with no pair of workers working together twice. Even after all possible pairs have worked once, all the workers together work for 11 more days to finish the work. Find the number of days in which all the workers together will finish the work?
Answer by jim_thompson5910(35256)

Answer: 2156 days
-----------------------------------------------------------------------------
-----------------------------------------------------------------------------

Explanation:

We have a pool of 66 workers to choose from.

We have two slots to fill: slot A, slot B

There are 66 choices for slot A.
There are 65 choices for slot B.
This makes 66*65 = 4290 permutations

Order doesn't matter so we divide this result by 2
4290/2 = 2145

There are 2145 different pairings possible.
So far, 2145 days have passed because each day is exactly one pair of workers.

Add on the 11 extra days
2145+11 = 2156

which is the final answer.

------------------------------------------------
Extra Info (optional)

If you wish to convert to years, then
2156/365 = 5.906849
so it takes roughly 5.906849 years

that's equivalent to roughly 5 years, 10.882188 months since
0.906849 years = (0.906849 yrs)*(12 mon/1 yr) = 10.882188 months

In short,
2156 days = 5.906849 years = 5 years, 10.882188 months
with the decimal values being approximate